a.An object moves in a straight line starting from rest. There are two stage in The Journey. 1. It gain speed uniformly for 2.0second and attains a 8.0ms-¹ 2.it continue at this speed for a further 1.5second. Sketch a graph of speed against time Find(1)the acceleration in stage (a) (2)the acceleration in stage (b) (3)the total distance moved during stages (a)and(b)
1
Expert's answer
2020-10-29T07:02:05-0400

(1) The acceleration in stage (a) is how quickly the velocity changes, or:


"a=\frac{v_f-v_i}{t_1}=\frac{8-0}{2}=4\text{ m/s}^2."

(2) The acceleration in stage (b) is zero because during 1.5 second the speed did not change.

(3) The entire distance is


"d=\frac{at_1^2}{2}+v_ft_2=\frac{4\cdot2^2}{2}+8\cdot1.5=20\text{ m}."

The speed-time graph is